OpenStudy (vshiroky):

Can someone please explain step by step... The furniture in a dollhouse was designed on a 1:30 scale of the original furniture. If an original bed measured 5 feet by 8 feet, what are the measurements of the bed in the dollhouse?

OpenStudy (johnweldon1993):

So this ratio means that the dollhouse furniture is 30 times smaller than the original furniture... So if the original was 5 by 8 The new will be \[\large \frac{5}{30} \text{by} \frac{8}{30}\]

OpenStudy (vshiroky):

What is my next step?

OpenStudy (johnweldon1993):

Just to divide them \[\large \frac{5}{30} = \space ?\] and \[\large \frac{8}{30} = \space ?\]

OpenStudy (vshiroky):

5 divided by 30?

OpenStudy (johnweldon1993):

Correct